Vocaloid:

find max and min of x^2 + y^2 subject to the constraint x^2 - 8x +y^2 - 10y = 0

Vocaloid:

min is pretty obvious but having trouble finding the maximum :/ I've used lagrange multipliers and end up with x = 0 and y = 10, giving 100 as the sol'n but this is incorrect

jhonyy9:

- one idea for this : for min. x=8 and y=10 we get x^2 +y^2 = 64 +100 = 164 this values i think bc. when x=8 and y=10 substituting these values inside this equation we get 0=0 for other different value